A screen is placed 1.70 m behind a single slit. The central maximum in the resulting diffraction pattern on the screen is 1.40 cm wide-that is, the two first-order diffraction minima are separated by 1.40 cm.

What is the distance between the two second-order minima?

Express your answer to three significant figures.
Answer #1

fringe width for 1st order minima

x = n L D/d

since fringe width is directly proportional to n

so

x2/ x1 = n2/n1

x2 = (n2/n1)x2

for 2nd order

x2' = 2 x2 = 2* ( 2/1) * (1.4 /2)

x2' = 2.8 cm

======